# Learning Equilibria from Data: Provably Efficient Multi-Agent Imitation Learning

### v1

@@ Appendix I Extension to $n$-player general-sum Games
Next, we want to sketch the extension of Algorithm 2 for $n$-player general-sum Game. We only give the extension for this algorithm as the ideas translate analogously to Algorithm 1. In a first step we have to adjust the decomposition in (5). We get
$\left({\frac{1}{K}\sum^{K}_{k=1}\sum^{n}_{i=1}\max_{\pi_{i}\in\Pi}\left\langle{d_{0}},{V^{\pi_{i},\pi^{k}_{-i}}}\right\rangle-\left\langle{d_{0}},{V^{\pi^{k}_{i},\pi^{k}_{-i}}}\right\rangle}\right)^{2}$
$=\left({\frac{1}{K}\sum^{K}_{k=1}\sum^{n}_{i=1}\left\langle{d_{0}},{V^{\pi^{\star,k}_{i},\pi^{k}_{-i}}-V^{\pi^{k}_{i},\pi^{k}_{-i}}}\right\rangle}\right)^{2}$
$=\left({\frac{1}{K}\sum^{K}_{k=1}\sum^{n}_{i=1}\left\langle{d_{0}},{V^{\pi^{\star,k}_{i},\pi^{k}_{-i}}-V^{\pi_{E_{i}},\pi_{E_{-i}}}+V^{\pi_{E_{i}},\pi_{E_{-i}}}-V^{\pi^{k}_{i},\pi^{k}_{-i}}}\right\rangle}\right)^{2}$
$\leq\left({\frac{1}{K}\sum^{K}_{k=1}\sum^{n}_{i=1}\left\langle{d_{0}},{V^{\pi^{\star,k}_{i},\pi^{k}_{-i}}-V^{\pi^{\star,k}_{i},\pi_{E_{-i}}}+V^{\pi_{E_{i}},\pi_{E_{-i}}}-V^{\pi^{k}_{i},\pi^{k}_{-i}}}\right\rangle}\right)^{2}$
$\leq 2\left({\underbrace{\frac{1}{K}\sum^{K}_{k=1}\sum^{n}_{i=1}\left\langle{d_{0}},{V^{\pi^{\star,k}_{i},\pi^{k}_{-i}}-V^{\pi^{\star,k}_{i},\pi_{E_{-i}}}}\right\rangle}_{\mathrm{(i)Exploit-Gap}}}\right)^{2}$
$\quad+2\left({\underbrace{\frac{1}{K}\sum^{K}_{k=1}\sum^{n}_{i=1}\left\langle{d_{0}},{V^{\pi_{E_{i}},\pi_{E_{-i}}}-V^{\pi^{k}_{i},\pi^{k}_{-i}}}\right\rangle}_{\mathrm{(ii)Value-Gap}}}\right)^{2}$
$\leq 2n\left(\left({\frac{1}{K}\sum^{K}_{k=1}\left\langle{d_{0}},{V^{\mathrm{br}(\pi^{k}_{-1}),\pi^{k}_{-1}}-V^{\mathrm{br}(\pi^{k}_{-1}),\pi_{E_{-1}}}}\right\rangle}\right)^{2}\right.$
$\left.\quad+\ldots+\left({\frac{1}{K}\sum^{K}_{k=1}\left\langle{d_{0}},{V^{\mathrm{br}(\pi^{k}_{-n}),\pi^{k}_{-n}}-V^{\mathrm{br}(\pi^{k}_{-n}),\pi_{E_{-n}}}}\right\rangle}\right)^{2}\right)$
$\quad+2\left({\frac{1}{K}\sum^{K}_{k=1}\sum^{n}_{i=1}\left\langle{d_{0}},{V^{\pi_{E_{i}},\pi_{E_{-i}}}-V^{\pi^{k}_{i},\pi^{k}_{-i}}}\right\rangle}\right)^{2},$
where $\pi^{\star,k}_{i}\in\mathrm{br}(\pi^{k}_{-i}).$ We will focus on $\mathrm{(i)}$ from now on, as the Value Gap is easy to bound by eg the Total Variation as also done in (14) and discussed in  Tang et al. [2024]. In particular, we will focus on the composition for any player $i\in\{1,\ldots,n\}.$ For $(i)$, we cannot continue directly as done in proof of Theorem 4.2 as now the policies inside differ in $n-1$ other policies and therefore we cannot directly apply the performance difference lemma. Instead, we first have to do the following
$\mathrm{(i)}$ $=\left({\frac{1}{K}\sum^{K}_{k=1}\left\langle{d_{0}},{V^{\pi^{\star,k}_{i},\pi^{k}_{-i}}-V^{\pi^{\star,k}_{i},\pi_{E_{-i}}}}\right\rangle}\right)^{2}$
$=\left({\frac{1}{K}\sum^{K}_{k=1}\left\langle{d_{0}},{V^{\pi^{\star,k}_{i},(\pi^{k}_{1},\ldots,\pi^{k}_{i-1},\pi^{k}_{i+1},\ldots\pi^{k}_{n})}-V^{\pi^{\star,k}_{i},(\pi_{E_{1}},\ldots,\pi_{E_{i-1}},\pi_{E_{i+1}},\ldots,\pi_{E_{n}})}}\right\rangle}\right)^{2}$
$=\bigg{(}\frac{1}{K}\sum^{K}_{k=1}\left\langle d_{0},V^{\pi^{\star,k}_{i},(\pi^{k}_{1},\ldots,\pi^{k}_{i-1},\pi^{k}_{i+1},\ldots\pi^{k}_{n})}-V^{\pi^{\star,k}_{i},(\pi^{k}_{1},\ldots,\pi^{k}_{i-1},\pi^{k}_{i+1},\ldots,\pi^{k}_{n-1},\pi_{E_{n}})}\right.$
$\left.\qquad\qquad+V^{\pi^{\star,k}_{i},(\pi^{k}_{1},\ldots,\pi^{k}_{i-1},\pi^{k}_{i+1},\pi^{k}_{n-1},\ldots\pi_{E_{n}})}-V^{\pi^{\star,k}_{i},(\pi^{k}_{1},\ldots,\pi^{k}_{i-1},\pi^{k}_{i+1},\ldots,\pi_{E_{n-1}},\pi_{E_{n}})}\right.$
[…]
$\left.\qquad\qquad+V^{\pi^{\star,k}_{i},(\pi^{k}_{1},\ldots,\pi_{E_{i-1}},\pi_{E_{i+1}},\ldots,\pi_{E_{n}})}-V^{\pi^{\star,k}_{i},(\pi_{E_{1}},\ldots,\pi_{E_{i-1}},\pi_{E_{i+1}},\ldots,\pi_{E_{n}})}\right\rangle\bigg{)}^{2}$
$\leq(n-1)\left(\left(\frac{1}{K}\sum^{K}_{k=1}\left\langle d_{0},V^{\pi^{\star,k}_{i},(\pi^{k}_{1},\ldots,\pi^{k}_{i-1},\pi^{k}_{i+1},\ldots\pi^{k}_{n})}-V^{\pi^{\star,k}_{i},(\pi^{k}_{1},\ldots,\pi^{k}_{i-1},\pi^{k}_{i+1},\ldots,\pi^{k}_{n-1},\pi_{E_{n}})}\right\rangle\right)^{2}\right.$
[…]
$\left.+\left(\frac{1}{K}\sum^{K}_{k=1}\left\langle d_{0},V^{\pi^{\star,k}_{i},(\pi^{k}_{1},\ldots,\pi_{E_{i-1}},\pi_{E_{i+1}},\ldots,\pi_{E_{n}})}-V^{\pi^{\star,k}_{i},(\pi_{E_{1}},\ldots,\pi_{E_{i-1}},\pi_{E_{i+1}},\ldots,\pi_{E_{n}})}\right\rangle\right)^{2}\right)$
Note that by the telescopic sum construction, we now have that each difference of value function only differs in one policy and last we applied $(a+b)^{2}\leq 2(a^{2}+b^{2})$ We will now focus on one term for the exploit Gap for any $i$. Therefore, we can proceed similar as in proof Theorem 4.2 and by dividing out $K^{2}$, applying the performance difference lemma ($(n-1)$ times, for every player but player $i$), Cauchy Schwartz and Jensen we get
$(n-1)\left(\left(\frac{1}{K}\sum^{K}_{k=1}\left\langle d_{0},V^{\pi^{\star,k}_{i},(\pi^{k}_{1},\ldots,\pi^{k}_{i-1},\pi^{k}_{i+1},\ldots\pi^{k}_{n})}-V^{\pi^{\star,k}_{i},(\pi^{k}_{1},\ldots,\pi^{k}_{i-1},\pi^{k}_{i+1},\ldots,\pi^{k}_{n-1},\pi_{E_{n}})}\right\rangle\right)^{2}\right.$
[…]
$\left.+\left(\frac{1}{K}\sum^{K}_{k=1}\left\langle d_{0},V^{\pi^{\star,k}_{i},(\pi^{k}_{1},\ldots,\pi_{E_{i-1}},\pi_{E_{i+1}},\ldots,\pi_{E_{n}})}-V^{\pi^{\star,k}_{i},(\pi_{E_{1}},\ldots,\pi_{E_{i-1}},\pi_{E_{i+1}},\ldots,\pi_{E_{n}})}\right\rangle\right)^{2}\right)$
$\leq\frac{(n-1)\left|{{\mathcal{A}}_{\max}}\right|}{(1-\gamma)^{2}K}\left(\sum^{K}_{k=1}\mathbb{E}_{s\sim d^{\pi^{\star,k}_{i},\pi^{k}_{-i}}}\left[\left\|\pi^{k}_{n}(\cdot\mid s)-\pi_{E_{n}}(\cdot\mid s)\right\|^{2}\right]\right.$
[…]
$\left.\qquad\qquad\qquad\qquad+\sum^{K}_{k=1}\mathbb{E}_{s\sim d^{\pi^{\star,k}_{i},(\pi^{k}_{1},\pi_{E_{-(i,1)}})}}\left[\left\|\pi^{k}_{1}(\cdot\mid s)-\pi_{E_{1}}(\cdot\mid s)\right\|^{2}\right]\right)$
@@
Now, we have to run $n-1$ RL-inner loops. Following the same analysis as done in proof of Theorem 4.2, we get a total bound in the order of ${\mathcal{O}}(\frac{n^{2}\left|{{\mathcal{S}}}\right|^{4}\left|{{\mathcal{A}}_{\max}}\right|^{5}}{(1-\gamma)^{12}\epsilon^{8}})$. Similar steps can also be done for Algorithm 1 to obtain ${\mathcal{O}}(\frac{n^{2}\left|{{\mathcal{S}}}\right|\left|{{\mathcal{A}}_{\max}}\right|^{2}}{(1-\gamma)^{4}\epsilon^{4}})$. This shows that the algorithm design indeed allows to avoid the curse of multi-agents and instead scales polynomial in the number of agents $n$.
Appendix J Comparison to Lower Bound in Tang et al.
@@
In this section, we compare our result Theorem 3.2 to Theorem 4.3 in Tang et al. [2024]. In particular, we emphasize how their construction allows to avoid a linear regret in the case of a fully known transition model. Note that they consider a finite horizon setting and general-sum games with a correlated equilibrium expert. For a better readability we first restate their Theorem in a infinite horizon setting.
Theorem J.1 (Theorem 4.3 in Tang et al. [2024]).
@@
The Markov Game that they construct is given in Fig. 3.
@@
$P(\cdot|s_{0},a,b)=\begin{dcases*}s_{1}&\text{if $(a,b)=a_{2}a_{1}$,}\\ s_{4}&\text{otherwise.}\end{dcases*}$
@@
It follows immediately, that an expert with a Nash-gap of 00, i.e. an NE is given by the following policy pair
@@
and in the other states any action $a\in{\mathcal{A}}$ can be chosen. As the expert data is not covering data for state $s_{1}$, it only covers the blue path in Fig. 3, Tang et al. [2024] argue that any policy can be chosen for the learner in state $s_{1}$ and choosing $\mu(a_{1}\mid s_{1})=\nu(a_{1}\mid s_{1})=1.$ However, if we know the transition model, the learner can be steered to choose a robust action such that the learner will be taken back to states known from the expert data, i.e. state $s_{4}$, even when one agent would deviate from the current policy. This is highlighted in green in Fig. 3. The only action that is considered robust is action $a_{3}$ for both agents, exactly the action chosen from the expert. This implies that the learning policy in the case of a fully known transition model has
@@
This in contrast to our construction in Fig. 1, where even under a known transition there is no way to steer the learner to known paths. Therefore, Theorem 3.2 shows the necessity of the single agent deviation coefficient and separates MAIL form SAIL, where effective learning is possible under known transitions Rajaraman et al. [2020]. Additionally, we consider a Zero-Sum Markov Game, not considered in Tang et al. Tang et al. [2024].
Appendix K Experiments

### v2

@@ Appendix H Extension to $n$-player general-sum games
$\textrm{Nash-Gap}(\boldsymbol{\pi})\leq\frac{8n}{(1-\gamma)^{2}}\max_{i}\max_{\pi_{i}\in\mathrm{BR}(\pi_{-i})}\left\|\frac{d^{\pi_{i},\pi_{E_{-i}}}}{d^{\pi_{E_{i}},\pi_{E_{-i}}}}\right\|_{\infty}\sqrt{\frac{\left|{{\mathcal{S}}}\right|\left(\sum_{i}{\mathcal{A}}_{i}\right)\log^{2}(n\left|{{\mathcal{S}}}\right|/\delta)}{N}}$
Next, we want to sketch the extension of Algorithm˜2 for $n$-player general-sum games. We only give the extension for this algorithm as the ideas translate analogously to Algorithm˜1. In a first step we have to adjust the decomposition in (5). We get
$\left({\frac{1}{K}\sum^{K}_{k=1}\sum^{n}_{i=1}\max_{\pi_{i}\in\Pi}\left\langle{d_{0}},{V_{i}^{\pi_{i},\pi^{k}_{-i}}}\right\rangle-\left\langle{d_{0}},{V_{i}^{\pi^{k}_{i},\pi^{k}_{-i}}}\right\rangle}\right)^{2}$
$=\left({\frac{1}{K}\sum^{K}_{k=1}\sum^{n}_{i=1}\left\langle{d_{0}},{V_{i}^{\pi^{\star,k}_{i},\pi^{k}_{-i}}-V_{i}^{\pi^{k}_{i},\pi^{k}_{-i}}}\right\rangle}\right)^{2}$
$=\left({\frac{1}{K}\sum^{K}_{k=1}\sum^{n}_{i=1}\left\langle{d_{0}},{V_{i}^{\pi^{\star,k}_{i},\pi^{k}_{-i}}-V_{i}^{\pi_{E_{i}},\pi_{E_{-i}}}+V_{i}^{\pi_{E_{i}},\pi_{E_{-i}}}-V_{i}^{\pi^{k}_{i},\pi^{k}_{-i}}}\right\rangle}\right)^{2}$
$\leq\left({\frac{1}{K}\sum^{K}_{k=1}\sum^{n}_{i=1}\left\langle{d_{0}},{V_{i}^{\pi^{\star,k}_{i},\pi^{k}_{-i}}-V_{i}^{\pi^{\star,k}_{i},\pi_{E_{-i}}}+V_{i}^{\pi_{E_{i}},\pi_{E_{-i}}}-V_{i}^{\pi^{k}_{i},\pi^{k}_{-i}}}\right\rangle}\right)^{2}$
$\leq 2\left({\underbrace{\frac{1}{K}\sum^{K}_{k=1}\sum^{n}_{i=1}\left\langle{d_{0}},{V_{i}^{\pi^{\star,k}_{i},\pi^{k}_{-i}}-V_{i}^{\pi^{\star,k}_{i},\pi_{E_{-i}}}}\right\rangle}_{\mathrm{(i)Exploit-Gap}}}\right)^{2}$
$\quad+2\left({\underbrace{\frac{1}{K}\sum^{K}_{k=1}\sum^{n}_{i=1}\left\langle{d_{0}},{V_{i}^{\pi_{E_{i}},\pi_{E_{-i}}}-V_{i}^{\pi^{k}_{i},\pi^{k}_{-i}}}\right\rangle}_{\mathrm{(ii)Value-Gap}}}\right)^{2}$
$\leq 2n\left(\left({\frac{1}{K}\sum^{K}_{k=1}\left\langle{d_{0}},{V_{1}^{\mathrm{br}(\pi^{k}_{-1}),\pi^{k}_{-1}}-V_{1}^{\mathrm{br}(\pi^{k}_{-1}),\pi_{E_{-1}}}}\right\rangle}\right)^{2}\right$.
$\left.\quad+\ldots+\left({\frac{1}{K}\sum^{K}_{k=1}\left\langle{d_{0}},{V_{n}^{\mathrm{br}(\pi^{k}_{-n}),\pi^{k}_{-n}}-V_{n}^{\mathrm{br}(\pi^{k}_{-n}),\pi_{E_{-n}}}}\right\rangle}\right)^{2}\right)$
$\quad+2\left({\frac{1}{K}\sum^{K}_{k=1}\sum^{n}_{i=1}\left\langle{d_{0}},{V_{i}^{\pi_{E_{i}},\pi_{E_{-i}}}-V_{i}^{\pi^{k}_{i},\pi^{k}_{-i}}}\right\rangle}\right)^{2},$
where $\pi^{\star,k}_{i}\in\mathrm{br}(\pi^{k}_{-i}).$ We will focus on $\mathrm{(i)}$ first. Then, we tackle the Value Gap. In particular, we will focus on the composition for any player $i\in\{1,\ldots,n\}.$ For $(i)$, we cannot continue directly as done in proof of Theorem˜4.2 as now the policies inside differ in n1n-1 other policies and therefore we cannot directly apply the performance difference lemma. Instead, we first have to do the following
$\mathrm{(i)}$ $=\left({\frac{1}{K}\sum^{K}_{k=1}\left\langle{d_{0}},{V_{i}^{\pi^{\star,k}_{i},\pi^{k}_{-i}}-V_{i}^{\pi^{\star,k}_{i},\pi_{E_{-i}}}}\right\rangle}\right)^{2}$
$=\left({\frac{1}{K}\sum^{K}_{k=1}\left\langle{d_{0}},{V_{i}^{\pi^{\star,k}_{i},(\pi^{k}_{1},\ldots,\pi^{k}_{i-1},\pi^{k}_{i+1},\ldots\pi^{k}_{n})}-V_{i}^{\pi^{\star,k}_{i},(\pi_{E_{1}},\ldots,\pi_{E_{i-1}},\pi_{E_{i+1}},\ldots,\pi_{E_{n}})}}\right\rangle}\right)^{2}$
$=\bigg(\frac{1}{K}\sum^{K}_{k=1}\left\langle d_{0},V_{i}^{\pi^{\star,k}_{i},(\pi^{k}_{1},\ldots,\pi^{k}_{i-1},\pi^{k}_{i+1},\ldots\pi^{k}_{n})}-V_{i}^{\pi^{\star,k}_{i},(\pi^{k}_{1},\ldots,\pi^{k}_{i-1},\pi^{k}_{i+1},\ldots,\pi^{k}_{n-1},\pi_{E_{n}})}\right$.
$\left.\qquad\qquad+V_{i}^{\pi^{\star,k}_{i},(\pi^{k}_{1},\ldots,\pi^{k}_{i-1},\pi^{k}_{i+1},\pi^{k}_{n-1},\ldots\pi_{E_{n}})}-V_{i}^{\pi^{\star,k}_{i},(\pi^{k}_{1},\ldots,\pi^{k}_{i-1},\pi^{k}_{i+1},\ldots,\pi_{E_{n-1}},\pi_{E_{n}})}\right$.
$\qquad\qquad\quad\vdots$
$\left.\qquad\qquad+V_{i}^{\pi^{\star,k}_{i},(\pi^{k}_{1},\ldots,\pi_{E_{i-1}},\pi_{E_{i+1}},\ldots,\pi_{E_{n}})}-V_{i}^{\pi^{\star,k}_{i},(\pi_{E_{1}},\ldots,\pi_{E_{i-1}},\pi_{E_{i+1}},\ldots,\pi_{E_{n}})}\right\rangle\bigg)^{2}$
$\leq(n-1)\left(\left(\frac{1}{K}\sum^{K}_{k=1}\left\langle d_{0},V_{i}^{\pi^{\star,k}_{i},(\pi^{k}_{1},\ldots,\pi^{k}_{i-1},\pi^{k}_{i+1},\ldots\pi^{k}_{n})}-V_{i}^{\pi^{\star,k}_{i},(\pi^{k}_{1},\ldots,\pi^{k}_{i-1},\pi^{k}_{i+1},\ldots,\pi^{k}_{n-1},\pi_{E_{n}})}\right\rangle\right)^{2}\right$.
$\qquad\qquad\quad\vdots$
$\left.+\left(\frac{1}{K}\sum^{K}_{k=1}\left\langle d_{0},V_{i}^{\pi^{\star,k}_{i},(\pi^{k}_{1},\ldots,\pi_{E_{i-1}},\pi_{E_{i+1}},\ldots,\pi_{E_{n}})}-V_{i}^{\pi^{\star,k}_{i},(\pi_{E_{1}},\ldots,\pi_{E_{i-1}},\pi_{E_{i+1}},\ldots,\pi_{E_{n}})}\right\rangle\right)^{2}\right)$
Note that by the telescopic sum construction, we now have that each difference of value function only differs in one policy and last we applied $(a+b)^{2}\leq 2(a^{2}+b^{2})$ . We will now focus on one term for the exploit Gap for any $i$. Therefore, we can proceed similar as in proof Theorem˜4.2 and by dividing out $K^{2}$, applying the performance difference lemma ((n1)(n-1) times, for every player but player $i$), Cauchy Schwarz and Jensen we get
$(n-1)\left(\left(\frac{1}{K}\sum^{K}_{k=1}\left\langle d_{0},V_{i}^{\pi^{\star,k}_{i},(\pi^{k}_{1},\ldots,\pi^{k}_{i-1},\pi^{k}_{i+1},\ldots\pi^{k}_{n})}-V_{i}^{\pi^{\star,k}_{i},(\pi^{k}_{1},\ldots,\pi^{k}_{i-1},\pi^{k}_{i+1},\ldots,\pi^{k}_{n-1},\pi_{E_{n}})}\right\rangle\right)^{2}\right$.
$\quad\vdots$
$\left.+\left(\frac{1}{K}\sum^{K}_{k=1}\left\langle d_{0},V_{i}^{\pi^{\star,k}_{i},(\pi^{k}_{1},\ldots,\pi_{E_{i-1}},\pi_{E_{i+1}},\ldots,\pi_{E_{n}})}-V_{i}^{\pi^{\star,k}_{i},(\pi_{E_{1}},\ldots,\pi_{E_{i-1}},\pi_{E_{i+1}},\ldots,\pi_{E_{n}})}\right\rangle\right)^{2}\right)$
$\leq\frac{(n-1)\left|{A_{\max}}\right|}{(1-\gamma)^{2}K}\left(\sum^{K}_{k=1}\mathbb{E}_{s\sim d^{y^{\star,k}_{i,1},\pi^{k}_{-i}}}\left[\left\|\pi^{k}_{n}(\cdot\mid s)-\pi_{E_{n}}(\cdot\mid s)\right\|^{2}\right]+\varepsilon^{i,1}_{\mathrm{opt}}\right$.
$\qquad\qquad\qquad\qquad\qquad\vdots$
$\left.\qquad\qquad\qquad\qquad+\sum^{K}_{k=1}\mathbb{E}_{s\sim d^{y^{\star,k}_{i,n},(\pi^{k}_{1},\pi_{E_{-(i,1)}})}}\left[\left\|\pi^{k}_{1}(\cdot\mid s)-\pi_{E_{1}}(\cdot\mid s)\right\|^{2}\right]+\varepsilon^{i,n}_{\mathrm{opt}}\right)$
$=\frac{(n-1)\left|{A_{\max}}\right|}{(1-\gamma)^{2}K}\sum^{K}_{k=1}\sum^{n}_{j\neq i}\mathbb{E}_{s\sim d^{y^{\star,k}_{ij},(\pi^{k}_{j:n}\pi_{\operatorname{E},1:j})_{-i}}}\left[{\left\|{\pi^{k}_{n-j+1}(s)-\pi_{\operatorname{E},n-j+1}(s)}\right\|^{2}}\right]+\varepsilon^{i,j}_{\mathrm{opt}}$
where we introduced n1n-1 RL inner loops for the player $i$ to approximate the maximum uncertainty response policy denoted by $y^{\star,k}_{ij}$ for $j\in[n]$. More precisely, we have
$\mathbb{E}_{s\sim d^{\pi^{\star,k}_{i},(\pi^{k}_{j:n}\pi_{\operatorname{E},1:j})_{-i}}}\left[{\left\|{\pi^{k}_{n-j+1}(s)-\pi_{\operatorname{E},n-j+1}(s)}\right\|^{2}}\right]$
$\leq\mathbb{E}_{s\sim d^{y^{\star,k}_{ij},(\pi^{k}_{j:n}\pi_{\operatorname{E},1:j})_{-i}}}\left[{\left\|{\pi^{k}_{n-j+1}(s)-\pi_{\operatorname{E},n-j+1}(s)}\right\|^{2}}\right]+\varepsilon^{ij}_{\mathrm{opt}}$
For the Value Gap, we obtain the following:
$\left({\frac{1}{K}\sum^{K}_{k=1}\sum^{n}_{i=1}\left\langle{d_{0}},{V_{i}^{\pi_{E_{i}},\pi_{E_{-i}}}-V_{i}^{\pi^{k}_{i},\pi^{k}_{-i}}}\right\rangle}\right)^{2}$ (15)
$\leq\left({\frac{2n}{1-\gamma}\frac{1}{K}\sum^{K}_{k=1}\mathbb{E}_{s\sim d^{\pi^{k}}}\left[\mathrm{TV}\left(\pi_{E}(\cdot\mid s),\pi^{k}(\cdot\mid s)\right)\right]}\right)^{2}$ (16)
$\leq\frac{4n^{2}}{(1-\gamma)^{2}}\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{s\sim d^{\pi_{k}}}\left[\mathrm{TV}^{2}\left(\pi_{E}(\cdot\mid s),\pi^{k}(\cdot\mid s)\right)\right]$ (17)
$=\frac{4n^{2}}{(1-\gamma)^{2}}\frac{1}{K}\sum_{k=1}^{K}\sum_{i=1}^{n}\mathbb{E}_{s\sim d^{\pi^{k}}}\left[\mathrm{TV}^{2}\left(\pi_{E_{i}}(\cdot\mid s),\pi_{i}^{k}(\cdot\mid s)\right)\right]$ (18)
where for the first inequality we have applied the performance difference lemma to the joint policies $\pi^{E}$ and $\pi^{k}$ as well as Hölder with $\left\|\cdot\right\|_{1}$ and $\left\|\cdot\right\|_{\infty}$ and bound the value function with its maximum value $\frac{1}{1-\gamma}.$ For the second inequality we used Cauchy Schwarz/Jensen’s inequality and in the last step that the policies are all conditional independent on $s$. Therefore, the whole objective can be upper bounded by
$\frac{4n^{2}}{(1-\gamma)^{2}}\frac{1}{K}\sum_{k=1}^{K}\sum_{i=1}^{n}\mathbb{E}_{s\sim d^{\pi^{k}}}\left[\mathrm{TV}^{2}\left(\pi_{E_{i}}(\cdot\mid s),\pi_{i}^{k}(\cdot\mid s)\right)\right]$
$+\frac{(n-1)\left|{A_{\max}}\right|}{(1-\gamma)^{2}K}\sum^{K}_{k=1}\sum^{n}_{i=1}\sum^{n}_{j\neq i}\mathbb{E}_{s\sim d^{y^{\star,k}_{ij},(\pi^{k}_{j:n}\pi_{\operatorname{E},1:j})_{-i}}}\left[{\left\|{\pi^{k}_{n-j+1}(s)-\pi_{\operatorname{E},n-j+1}(s)}\right\|^{2}}\right]+\varepsilon^{i,j}_{\mathrm{opt}}$
$=\frac{1}{K}\sum_{k=1}^{K}\sum_{i=1}^{n}\left(\frac{4n^{2}}{(1-\gamma)^{2}}\mathbb{E}_{s\sim d^{\pi^{k}}}\left[\mathrm{TV}^{2}\left(\pi_{E_{i}}(\cdot\mid s),\pi_{i}^{k}(\cdot\mid s)\right)\right]\right$.
$\left.+\frac{(n-1)\left|{A_{\max}}\right|}{(1-\gamma)^{2}}\sum^{n}_{j\neq i}\mathbb{E}_{s\sim d^{y^{\star,k}_{ij},(\pi^{k}_{j:n}\pi_{\operatorname{E},1:j})_{-i}}}\left[{\left\|{\pi^{k}_{n-j+1}(s)-\pi_{\operatorname{E},n-j+1}(s)}\right\|^{2}}\right]+\varepsilon^{i,j}_{\mathrm{opt}}\right)$
$\leq\frac{n^{2}\left|{A_{\max}}\right|}{(1-\gamma)^{2}}\frac{1}{K}\sum_{k=1}^{K}\sum_{i=1}^{n}\left(\mathbb{E}_{s\sim d^{\pi^{k}}}\left[\left\|{\pi_{E_{i}}(s)-\pi_{i}^{k}(s)}\right\|^{2}\right]\right$.
$\left.+\sum^{n}_{j\neq i}\mathbb{E}_{s\sim d^{y^{\star,k}_{ij},(\pi^{k}_{j:n}\pi_{\operatorname{E},1:j})_{-i}}}\left[{\left\|{\pi^{k}_{n-j+1}(s)-\pi_{\operatorname{E},n-j+1}(s)}\right\|^{2}}\right]+\varepsilon^{i,j}_{\mathrm{opt}}\right),$
where we used that $\mathrm{TV}^{2}(\pi_{E_{i}},\pi_{i})\leq\frac{A_{i}}{4}\left\|{\pi_{E_{i}}-\pi_{i}}\right\|^{2}.$
Now note that we can sample according to all state occupancy measures. In particular, we have $S^{i,j}_{k}\sim d^{y^{\star,k}_{ij},(\pi^{k}_{j:n}\pi_{\operatorname{E},1:j})_{-i}}$ and $S_{k}^{\pi_{k}}\sim d^{\pi^{k}}.$ Again adding and subtracting $\left\|{\pi_{E_{i}}(\cdot\mid S^{\pi^{k}}_{k})-\pi^{k}(\cdot\mid S^{\pi^{k}}_{k})}\right\|^{2}$ and $\left\|{\pi_{E_{i}}(\cdot\mid S^{i,j}_{k})-\pi^{k}(\cdot\mid S^{i,j}_{k})}\right\|^{2}$ we get
$\frac{n^{2}\left|{A_{\max}}\right|}{(1-\gamma)^{2}}\frac{1}{K}\sum_{i=k}^{K}\sum_{i=1}^{n}$
$\left(\sum^{n}_{j\neq i}\mathbb{E}_{s\sim d^{y^{\star,k}_{ij},(\pi^{k}_{j:n}\pi_{\operatorname{E},1:j})_{-i}}}\left[{\left\|{\pi^{k}_{n-j+1}(\cdot\mid s)-\pi_{\operatorname{E},n-j+1}(\cdot\mid s)}\right\|^{2}}\right]\right$.
$\left.\quad-\left\|{\pi_{E_{,}n-j+1}(\cdot\mid S^{i,j}_{k})-\pi_{n-j+1}^{k}(\cdot\mid S^{i,j}_{k})}\right\|^{2}\right$.
$\left.\quad+\mathbb{E}_{s\sim d^{\pi^{k}}}\left[\left\|{\pi_{E_{i}}(\cdot\mid s)-\pi_{i}^{k}(\cdot\mid s)}\right\|^{2}\right]-\left\|{\pi_{E_{i}}(\cdot\mid S^{\pi^{k}}_{k})-\pi^{k}(\cdot\mid S^{\pi^{k}}_{k})}\right\|^{2}\right$$\quad$ (Martingale)
$\left.\quad+\sum^{n}_{j\neq i}\left\|{\pi_{\operatorname{E},n-j+1}(\cdot\mid S^{i,j}_{k})-\pi_{n-j+1}^{k}(\cdot\mid S^{i,j}_{k})}\right\|^{2}\right$.
$\left.\quad+\left\|{\pi_{E_{i}}(\cdot\mid S^{\pi^{k}}_{k})-\pi^{k}(\cdot\mid S^{\pi^{k}}_{k})}\right\|^{2}\right)\quad$ (Regret)
$\quad+\frac{n^{3}\left|{A_{\max}}\right|}{(1-\gamma)^{2}}\varepsilon_{\mathrm{opt}}$
Note that now again, we have analogously to (12) a martingale term here and a regret term that we can control by updating the policies with online mirror descent. In particular, we can rearrange the regret term as follows
@@
At this point, following an analogous analysis as done in proof of Theorem˜4.2, we get a total bound in the order of ${\mathcal{O}}(\frac{\mathrm{poly}(n,\left|{{\mathcal{S}}}\right|,\left|{A_{\max}}\right|,(1-\gamma)^{-1})}{\varepsilon^{8}})$. Similar steps can also be done for Algorithm˜1 to obtain ${\mathcal{O}}(\frac{\mathrm{poly}(n,\left|{{\mathcal{S}}}\right|,\left|{A_{\max}}\right|,(1-\gamma)^{-1})}{\varepsilon^{4}})$. This shows that the algorithm design indeed allows to avoid the curse of multi-agents and instead scales polynomial in the number of agents $n$.
Appendix I Comparison to Lower Bound in Tang et al.
@@
In this section, we compare our result Theorem˜3.2 to Theorem 4.3 in Tang et al. [2024]. In particular, we emphasize how their construction allows to avoid a linear regret in the case of a fully known transition model. Note that they consider a finite horizon setting and general-sum games with a correlated equilibrium expert. For a better readability we first restate their Theorem in a infinite horizon setting.
Theorem I.1 (Theorem 4.3 in Tang et al. [2024]).
@@
The Markov Game that they construct is given in Fig.˜3.
@@
[…]
@@
It follows immediately, that an expert with a Nash-gap of 0, i.e. an NE is given by the following policy pair
@@
and in the other states any action $a\in{\mathcal{A}}$ can be chosen. As the expert data is not covering data for state $s_{1}$, it only covers the blue path in Fig.˜3, Tang et al. [2024] argue that any policy can be chosen for the learner in state $s_{1}$ and choosing $\mu(a_{1}\mid s_{1})=\nu(a_{1}\mid s_{1})=1.$ However, if we know the transition model, the learner can be steered to choose a robust action such that the learner will be taken back to states known from the expert data, i.e. state $s_{4}$, even when one agent would deviate from the current policy. This is highlighted in green in Fig.˜3. The only action that is considered robust is action $a_{3}$ for both agents, exactly the action chosen from the expert. This implies that the learning policy in the case of a fully known transition model has
@@
This in contrast to our construction in Fig.˜1, where even under a known transition there is no way to steer the learner to known paths. Therefore, Theorem˜3.2 shows the necessity of the expert agent deviation coefficient and separates MAIL form SAIL, where effective learning is possible under known transitions Rajaraman et al. [2020]. Additionally, we consider a Zero-Sum Markov Game, not considered in Tang et al. Tang et al. [2024].
Appendix J Experiments